OpenStudy (a1234):

Find the solutions of sin5x + sin3x = 0 in the interval [0, 2pi).

OpenStudy (a1234):

I got to 2sin4x cosx = 0 and got the solutions 0, pi, pi/2, and 3pi/2. I feel like some solutions are missing...

OpenStudy (tkhunny):

What did you do to obtain that answer set?

OpenStudy (a1234):

2sin4x = 0 cosx = 0 2sin4x = 0 equals sin4x = 0

OpenStudy (a1234):

@tkhunny

OpenStudy (tkhunny):

?? Where did you get sin(4x)?

OpenStudy (a1234):

Divided the 2 from both sides...not sure if it would work

OpenStudy (tkhunny):

Try this: \(\sin(5x) = \sin(4x + x) = \sin(4x)\cos(x) + \cos(4x)\sin(x)\) You do this one: \(\sin(3x) = \sin(4x - x) = \)